# MoE-CAP: Benchmarking Cost, Accuracy and Performance of Sparse Mixture-of-Experts Systems

### v5

@@ Abstract
The sparse Mixture-of-Experts (MoE) architecture is increasingly favored for scaling Large Language Models (LLMs) efficiently, but it depends on heterogeneous compute and memory resources. These factors jointly affect system Cost, Accuracy, and Performance (CAP), making trade-offs inevitable. Existing benchmarks often fail to capture these trade-offs accurately, complicating practical deployment decisions. To address this, we introduce MoE-CAP, a benchmark specifically designed for MoE systems. Our analysis reveals that achieving an optimal balance across CAP is difficult with current hardware; MoE systems typically optimize two of the three dimensions at the expense of the third—a dynamic we term the MoE-CAP trade-off. To visualize this, we propose the CAP Radar Diagram. We further introduce sparsity-aware performance metrics—Sparse Memory Bandwidth Utilization (S-MBU) and Sparse Model FLOPS Utilization (S-MFU)—to enable accurate performance benchmarking of MoE systems across diverse hardware platforms and deployment scenarios. Our code has been released at: https://github.com/sparse-generative-ai/MoE-CAP
@@ 1 Introduction
Recent large language models (LLMs) are increasingly adopting sparse Mixture-of-Experts (MoE) architectures, notable examples of which include Switch-C [14], DBRX, Mixtral-8x22B [25], Snowflake Arctic [39], Grok-1 [46], DeepSeek-MoE [10], and Qwen1.5-MoE [4]. These models utilize sparse experts grouped into an MoE layer, and these experts are selectively activated through a router (or a gating network). By routing tokens to a subset of experts, MoEs achieve sub-linear computational costs compared to their dense equivalents, which allows building trillion-parameter-scale LLMs.
Current MoE systems exhibit increasing complexity, driven by two main factors: (i) There is enhanced sophistication in the design of sparse MoE layers and gating networks (or routers), which differ in sparsity characteristics across various MoE models—we define sparsity in MoE systems as the ratio of activated to total parameters per token; (ii) MoEs demonstrate sub-linear computational complexity, enabling the offloading of less frequently activated experts onto external memory and processors. This approach reduces dependence on costly High Bandwidth Memory (HBM) on GPUs. Consequently, the complexity of servers hosting MoE systems has escalated, with these servers typically featuring heterogeneous compute, memory, and communication resources, arranged in a multi-tier architecture. For instance, modern MoE systems increasingly offload experts to DRAM and SSDs [49, 13] and delegate part of the computation to CPUs [26].
Practitioners of MoE systems are actively seeking methods to benchmark their cost, accuracy (in downstream tasks), and performance (time and memory efficiency) in order to optimize their deployment. However, this benchmarking is challenging for the following reasons: (i) Poor understanding of the relation between cost, accuracy and performance: real-world deployments of MoE systems frequently reveal underestimated costs, under-achievements in performance benefits, and compromised accuracy. Clear principles are needed to help practitioners effectively evaluate and understand the complex interplay between these factors, (ii) Inadequate system cost and performance assessment metrics: Existing metrics like Memory Bandwidth Utilization (MBU) [1] and Model FLOPS Utilization (MFU) [8] fail to account for the sparse activation patterns of experts in MoE systems. This oversight leads to overestimated memory and compute costs. Additionally, current benchmarks predominantly estimate costs based on GPU usage alone. However, modern MoE systems increasingly rely on heterogeneous processors and multi-tier memory resources. Ignoring these factors yields inaccurate cost estimations.
@@ 2 Background and Motivation
MoE systems exhibit sub-linear computational complexity but still require increased memory to accommodate all potentially activated experts. To enhance memory efficiency, several system designs have been explored recently: (i) Designs applying quantization to experts, such as unified quantization for all experts—e.g., GPTQ [16], AWQ [32], and SmoothQuant [47]—and expert-specific adaptive quantization—e.g., QMoE [15] and MoQE [27]. (ii) Designs offloading experts to external memory. Examples include MoE-layer-wise parameter offloading—e.g., DeepSpeed-Inference [2] and SwapAdvisor [23]—and fine-grained expert-level offloading—e.g., MoE-Infinity [49], Brainstorm [9], Mixtral-Offloading [13]. (iii) Designs utilizing CPU cores for sharing partial computation from GPUs, such as computing less input-intensive experts on CPUs— e.g., Fiddler [26].
@@ 2.1 Why existing benchmarks fall short with MoE systems
Many LLM benchmarks have been proposed over the past few years, including MLPerf [38], ML-Energy [51], Open-LLM-Leaderboard [5], LLM-Perf [24], TensorDock [40], and Artificial Analysis [3]. To evaluate LLM systems, these benchmarks primarily use three types of metrics: (i) overall system performance metrics, such as token throughput, prefilling time, and decoding time, (ii) cost metrics, like tokens-per-dollar and tokens-per-kilowatt, focusing solely on GPU usage, and (iii) fine-grained system performance metrics, such as MBU and MFU, to facilitate the understanding of memory and compute bottlenecks, as discussed in recent MoE articles [1]. However, when these benchmarks are applied to MoE systems, they demonstrate several limitations.
@@
(2) Inaccurate system cost and performance assessment metrics. The commonly used MBU [1] and MFU [8] metrics fail to consider the selective activation within sparse MoE layers. MBU measures memory bandwidth utilization as $\text{MBU}=\frac{B_{\text{achieved}}}{B_{\text{peak}}}=\frac{(S_{\text{model}}+S_{\text{KV}})/\text{TPOT}}{B_{\text{peak}}}$, where $S_{\text{model}}$ represents the full model size and $S_{\text{KV}}$ is the KV cache size, $B_{\text{achieved}}$ refers to the achieved memory bandwidth per forward pass, $B_{\text{peak}}$ is the peak bandwidth of the underlying hardware. Similarly, MFU measures compute utilization as $\textsc{MFU}=\left(T_{\text{token}}\times F_{\text{token}}\right)/F_{\text{peak}}$ , assuming all parameters participate in computation, where $T_{\text{token}}$ is the token throughput, $F_{\text{token}}$ represents the FLOPs required per token, and $F_{\text{peak}}$ indicates the peak FLOPs capacity of the hardware. However, as illustrated in Figure 2, these metrics significantly overestimate resource utilization by assuming all experts are active, especially when batch size >1. For example, when only one expert out of three is activated (case a), both MBU and MFU are inflated by 3×. Even with shared experts (case c), the metrics still overestimate utilization by 1.5×. This often leads operators to substantially over-provision GPUs, causing significant resource waste.
@@ 3 The CAP Benchmarking Method
    Accuracy (A): Accuracy in MoE systems is broadly defined to encompass a range of evaluation metrics commonly used in prominent LLM benchmarks, such as the Open LLM Leaderboard [5], which reflect real-world applications of large language models. To support comprehensive task evaluation—spanning multiple-choice, reasoning, and QA tasks—we adopt diverse metrics including exact match, F1 score, and win rate, ensuring broad coverage across various evaluation scenarios.
@@ 3.1 Complete cost model of MoE systems
We apply Equation (3) to combine both cost models and performance metrics. We aim to determine the per-token cost in units of dollars, denoted as $C_{\textrm{token}}$. The energy cost in kWh needs to be combined with the local energy price to estimate the cost, as it can vary between industrial/personal energy usage and by region [41]. The $C_{\textrm{token}}$ is calculated over the lifetime $R$ of the deployment and averaged across all generated tokens. To further explain our cost model, we illustrate the use cases of our cost model in Appendix A.2.
@@ 3.2 Understanding strengths of MoE systems through CAP analysis
MoE systems for Performance and Accuracy (PA). Improving performance without compromising accuracy can be achieved through two primary approaches: (i) scaling up device memory capacity by utilizing high-end GPUs with large memory (e.g., H200 NVL with 141 GB and MI300X with 192 GB), and (ii) scaling out memory capacity via parallel and distributed computing, leveraging technologies such as NVLinks, NVSwitch, and InfiniBand, along with techniques like Tensor, Pipeline, and Expert parallelism [30, 21, 31, 28, 37]. However, these approaches face significant drawbacks: the exponential increase in system costs due to the growing manufacturing complexity of high-speed memory devices and the sub-linear scaling efficiency of distributed systems, where communication increasingly becomes a bottleneck at larger scales.
MoE systems for Performance and Cost (PC). To enhance system performance without increasing costs, effective model compression techniques such as quantization and sparsification can be adopted for MoE models. Specifically, quantization reduces model size by mapping parameters to lower precisions, such as 8-bit [11, 50], 4-bit [32, 16], or even 2-bit [12, 7], and accelerates latency through advanced system implementations [17, 44] featuring optimized data layouts and efficient dequantization. Although MoE models inherently exhibit sparsity, their application in long-context scenarios faces challenges due to the quadratic complexity of attention mechanisms. Sparse attention, therefore, emerges as a compelling solution for MoE models [48, 19, 33, 52]. However, compression typically introduces accuracy degradation due to the limited representational capacity of compressed parameters.
MoE systems for Cost and Accuracy (CA). In constrained hardware environments, some systems aim to maintain model accuracy within a limited budget (e.g., relying on low-end GPUs or a restricted number of GPUs). Techniques such as memory swapping [49, 6] are employed to offload model parameters or KV caches based on sparsity. Additionally, some systems utilize CPUs to assist GPU operations during model inference or generation [26]. These systems however, still introduce performance degradation.
@@ 4.1 Sparse model bandwidth utilization
When designing a new sparsity-aware MBU, we want this new metric to be generally applicable to all types of MoE models we are aware of by far. These different sparsity patterns of MoE models can be summarized in Figure 2. From this figure, we can derive several requirements for the new MBU. First, the metric must capture the set of experts activated by a batch of input tokens. For example, in case (a), a sparse feed-forward (FF) layer contains three experts, and each token is routed to its top-1 expert. Both tokens are routed to the same expert, so the accessed memory corresponds to a single expert of size $S$. In contrast, case (b) routes the tokens to two distinct experts, doubling the accessed memory to $2S$. Second, the metrics must account for different routing mechanisms, such as shared experts introduced by recent MoE studies [10, 4] (illustrated in Figure 2 (c)).
@@
Besides, dense models are a special case of equation (4) with $n_{\text{expert}}=1$ and $\forall i,\mathbbm{1}[l,i]=1$. Therefore, our definition is also suitable for model architectures where not all layers are MoE layers, e.g., in Switch Transformers [14]. We further validate S-MBU accuracy on MoE models; detailed results are provided in Appendix A.4.
@@ 4.3 Benchmarking use cases of sparsity-aware performance metrics
Figure 6 illustrates the relationship between batch size and model sparsity, based on profiling using our custom benchmarking tool built on the latest vLLM release and evaluated on the MATH [20] dataset. In scenarios such as personal inference and post-training—characterized by small batch sizes (1–16)—all models exhibit high sparsity. At batch size 8, for instance, only 53.05%, 46.79%, and 18.44% of parameters are activated for DeepSeek-V2-Lite, Qwen1.5-MoE, and DeepSeek-R1, respectively. This significantly reduces bandwidth demands, making these models viable for deployment on cost-efficient hardware.
@@ 5 Benchmark Implementation
Dataset support. We evaluate all models on four representative benchmarks: MMLU, GSM8K, MATH, Arena-Hard, and LongBench. MMLU covers 57 diverse multiple-choice tasks to assess factuality and reasoning across domains. GSM8K and MATH focus on mathematical reasoning, with problems requiring multi-step solutions and short-form generation at varying difficulty levels. Arena-Hard evaluates long-form generation using 500 complex user queries from Chatbot Arena, judged by GPT-4-Turbo. It shows strong agreement with human preferences (89.1%) and better separability than other benchmarks. LongBench is a benchmark of long-context questions, making it well-suited for evaluating prefill-heavy workloads. Together, these benchmarks comprehensively test MoE LLMs across multiple-choice, short-form, and long-form generation tasks—covering knowledge, reasoning, and output quality—and are widely used in leading LLM leaderboards [5, 24, 22].
@@ A.1 MoE-CAP Limitation and Future Work
Toward real-world deployment scenarios. We also aim to evaluate MoE systems in diverse cloud deployment settings, including serverless model endpoints [18], elastic infrastructure [43, 45, 34], and spot-instance pricing environments [35, 36, 42]. This will enable more realistic assessments of cost-performance trade-offs under variable resource conditions.
@@ A.7 Practical bandwidth and compute requirement calculation
We calculated theoretical performance through existing tools like LLM-Analysis [29] and LLM-Viewer [53]. In practice, MoE systems often suffer inefficiencies due to redundant data transfers or calculations, resulting in performance losses. Consequently, the calculated theoretical bandwidth or OPS may not meet the required application performance (e.g., throughput). To determine the practical hardware requirements for a given set of inputs, it is essential to consider hardware utilization. Thus, the practical performance requirement is defined as:

### v6

@@ Abstract
The sparse Mixture-of-Experts (MoE) architecture is increasingly favored for scaling Large Language Models (LLMs) efficiently, but it depends on heterogeneous compute and memory resources. These factors jointly affect system Cost, Accuracy, and Performance (CAP), making trade-offs inevitable. Existing benchmarks often fail to capture these trade-offs accurately, complicating practical deployment decisions. To address this, we introduce MoE-CAP, a benchmark specifically designed for MoE systems. Our analysis reveals that achieving an optimal balance across CAP is difficult with current hardware; MoE systems typically optimize two of the three dimensions at the expense of the third—a dynamic we term the MoE-CAP trade-off. To visualize this, we propose the CAP Radar Diagram. We further introduce sparsity-aware performance metrics—Sparse Memory Bandwidth Utilization (S-MBU) and Sparse Model FLOPS Utilization (S-MFU)—to enable accurate performance benchmarking of MoE systems across diverse hardware platforms and deployment scenarios. Our code has been released at: https://github.com/Auto-CAP/MoE-CAP
@@ 1 Introduction
Recent large language models (LLMs) are increasingly adopting sparse Mixture-of-Experts (MoE) architectures, notable examples of which include Switch-C [switch], DBRX, Mixtral-8x22B [mixtral], Snowflake Arctic [snowflake-arctic], Grok-1 [grok-1], DeepSeek-MoE [deepseekmoe], and Qwen1.5-MoE [qwen]. These models utilize sparse experts grouped into an MoE layer, and these experts are selectively activated through a router (or a gating network). By routing tokens to a subset of experts, MoEs achieve sub-linear computational costs compared to their dense equivalents, which allows building trillion-parameter-scale LLMs.
Current MoE systems exhibit increasing complexity, driven by two main factors: (i) There is enhanced sophistication in the design of sparse MoE layers and gating networks (or routers), which differ in sparsity characteristics across various MoE models—we define sparsity in MoE systems as the ratio of activated to total parameters per token; (ii) MoEs demonstrate sub-linear computational complexity, enabling the offloading of less frequently activated experts onto external memory and processors. This approach reduces dependence on costly High Bandwidth Memory (HBM) on GPUs. Consequently, the complexity of servers hosting MoE systems has escalated, with these servers typically featuring heterogeneous compute, memory, and communication resources, arranged in a multi-tier architecture. For instance, modern MoE systems increasingly offload experts to DRAM and SSDs [xue2024moeinfinity, eliseev2023fast] and delegate part of the computation to CPUs [fiddler].
Practitioners of MoE systems are actively seeking methods to benchmark their cost, accuracy (in downstream tasks), and performance (time and memory efficiency) in order to optimize their deployment. However, this benchmarking is challenging for the following reasons: (i) Poor understanding of the relation between cost, accuracy and performance: real-world deployments of MoE systems frequently reveal underestimated costs, under-achievements in performance benefits, and compromised accuracy. Clear principles are needed to help practitioners effectively evaluate and understand the complex interplay between these factors, (ii) Inadequate system cost and performance assessment metrics: Existing metrics like Memory Bandwidth Utilization (MBU) [databricks-mbu] and Model FLOPS Utilization (MFU) [Chowdhery2022PaLMSL] fail to account for the sparse activation patterns of experts in MoE systems. This oversight leads to overestimated memory and compute costs. Additionally, current benchmarks predominantly estimate costs based on GPU usage alone. However, modern MoE systems increasingly rely on heterogeneous processors and multi-tier memory resources. Ignoring these factors yields inaccurate cost estimations.
@@ 2 Background and Motivation
MoE systems exhibit sub-linear computational complexity but still require increased memory to accommodate all potentially activated experts. To enhance memory efficiency, several system designs have been explored recently: (i) Designs applying quantization to experts, such as unified quantization for all experts—e.g., GPTQ [gptq], AWQ [awq], and SmoothQuant [smooth-quant]—and expert-specific adaptive quantization—e.g., QMoE [qmoe] and MoQE [moqe]. (ii) Designs offloading experts to external memory. Examples include MoE-layer-wise parameter offloading—e.g., DeepSpeed-Inference [deepspeed-inference] and SwapAdvisor [swapadvisor]—and fine-grained expert-level offloading—e.g., MoE-Infinity [xue2024moeinfinity], Brainstorm [brainstorm], Mixtral-Offloading [eliseev2023fast]. (iii) Designs utilizing CPU cores for sharing partial computation from GPUs, such as computing less input-intensive experts on CPUs— e.g., Fiddler [fiddler].
@@ 2.1 Why existing benchmarks fall short with MoE systems
Many LLM benchmarks have been proposed over the past few years, including MLPerf [mlperf], ML-Energy [mlenergy], Open-LLM-Leaderboard [open-llm-leaderboard], LLM-Perf [llm-perf-leaderboard], TensorDock [tensordock], and Artificial Analysis [artificialanalysis]. To evaluate LLM systems, these benchmarks primarily use three types of metrics: (i) overall system performance metrics, such as token throughput, prefilling time, and decoding time, (ii) cost metrics, like tokens-per-dollar and tokens-per-kilowatt, focusing solely on GPU usage, and (iii) fine-grained system performance metrics, such as MBU and MFU, to facilitate the understanding of memory and compute bottlenecks, as discussed in recent MoE articles [databricks-mbu]. However, when these benchmarks are applied to MoE systems, they demonstrate several limitations.
@@
(2) Inaccurate system cost and performance assessment metrics. The commonly used MBU [databricks-mbu] and MFU [Chowdhery2022PaLMSL] metrics fail to consider the selective activation within sparse MoE layers. MBU measures memory bandwidth utilization as $\text{MBU}=\frac{B_{\text{achieved}}}{B_{\text{peak}}}=\frac{(S_{\text{model}}+S_{\text{KV}})/\text{TPOT}}{B_{\text{peak}}}$, where $S_{\text{model}}$ represents the full model size and $S_{\text{KV}}$ is the KV cache size, $B_{\text{achieved}}$ refers to the achieved memory bandwidth per forward pass, $B_{\text{peak}}$ is the peak bandwidth of the underlying hardware. Similarly, MFU measures compute utilization as $\textsc{MFU}=\left(T_{\text{token}}\times F_{\text{token}}\right)/F_{\text{peak}}$ , assuming all parameters participate in computation, where $T_{\text{token}}$ is the token throughput, $F_{\text{token}}$ represents the FLOPs required per token, and $F_{\text{peak}}$ indicates the peak FLOPs capacity of the hardware. However, as illustrated in Figure 2, these metrics significantly overestimate resource utilization by assuming all experts are active, especially when batch size >1. For example, when only one expert out of three is activated (case a), both MBU and MFU are inflated by 3×. Even with shared experts (case c), the metrics still overestimate utilization by 1.5×. This often leads operators to substantially over-provision GPUs, causing significant resource waste.
@@ 3 The CAP Benchmarking Method
    Accuracy (A): Accuracy in MoE systems is broadly defined to encompass a range of evaluation metrics commonly used in prominent LLM benchmarks, such as the Open LLM Leaderboard [open-llm-leaderboard], which reflect real-world applications of large language models. To support comprehensive task evaluation—spanning multiple-choice, reasoning, and QA tasks—we adopt diverse metrics including exact match, F1 score, and win rate, ensuring broad coverage across various evaluation scenarios.
@@ 3.1 Complete cost model of MoE systems
We apply Equation (3) to combine both cost models and performance metrics. We aim to determine the per-token cost in units of dollars, denoted as $C_{\textrm{token}}$. The energy cost in kWh needs to be combined with the local energy price to estimate the cost, as it can vary between industrial/personal energy usage and by region [usenergyprice]. The $C_{\textrm{token}}$ is calculated over the lifetime $R$ of the deployment and averaged across all generated tokens. To further explain our cost model, we illustrate the use cases of our cost model in Appendix A.2.
@@ 3.2 Understanding strengths of MoE systems through CAP analysis
MoE systems for Performance and Accuracy (PA). Improving performance without compromising accuracy can be achieved through two primary approaches: (i) scaling up device memory capacity by utilizing high-end GPUs with large memory (e.g., H200 NVL with 141 GB and MI300X with 192 GB), and (ii) scaling out memory capacity via parallel and distributed computing, leveraging technologies such as NVLinks, NVSwitch, and InfiniBand, along with techniques like Tensor, Pipeline, and Expert parallelism [lina, fastgen, alpaserve, vllm, tensorrt-llm]. However, these approaches face significant drawbacks: the exponential increase in system costs due to the growing manufacturing complexity of high-speed memory devices and the sub-linear scaling efficiency of distributed systems, where communication increasingly becomes a bottleneck at larger scales.
MoE systems for Performance and Cost (PC). To enhance system performance without increasing costs, effective model compression techniques such as quantization and sparsification can be adopted for MoE models. Specifically, quantization reduces model size by mapping parameters to lower precisions, such as 8-bit [llm.int8, edge-moe], 4-bit [awq, gptq], or even 2-bit [du2024bitdistiller, efficientqat], and accelerates latency through advanced system implementations [frantar2024marlin, ladder-osdi24] featuring optimized data layouts and efficient dequantization. Although MoE models inherently exhibit sparsity, their application in long-context scenarios faces challenges due to the quadratic complexity of attention mechanisms. Sparse attention, therefore, emerges as a compelling solution for MoE models [xu2025xattention, gao2024seerattention, lu2025moba, nsa]. However, compression typically introduces accuracy degradation due to the limited representational capacity of compressed parameters.
MoE systems for Cost and Accuracy (CA). In constrained hardware environments, some systems aim to maintain model accuracy within a limited budget (e.g., relying on low-end GPUs or a restricted number of GPUs). Techniques such as memory swapping [xue2024moeinfinity, moe-lightning] are employed to offload model parameters or KV caches based on sparsity. Additionally, some systems utilize CPUs to assist GPU operations during model inference or generation [fiddler]. These systems however, still introduce performance degradation.
@@ 4.1 Sparse model bandwidth utilization
When designing a new sparsity-aware MBU, we want this new metric to be generally applicable to all types of MoE models we are aware of by far. These different sparsity patterns of MoE models can be summarized in Figure 2. From this figure, we can derive several requirements for the new MBU. First, the metric must capture the set of experts activated by a batch of input tokens. For example, in case (a), a sparse feed-forward (FF) layer contains three experts, and each token is routed to its top-1 expert. Both tokens are routed to the same expert, so the accessed memory corresponds to a single expert of size $S$. In contrast, case (b) routes the tokens to two distinct experts, doubling the accessed memory to $2S$. Second, the metrics must account for different routing mechanisms, such as shared experts introduced by recent MoE studies [deepseekmoe, qwen] (illustrated in Figure 2 (c)).
@@
Besides, dense models are a special case of equation (4) with $n_{\text{expert}}=1$ and $\forall i,\mathbbm{1}[l,i]=1$. Therefore, our definition is also suitable for model architectures where not all layers are MoE layers, e.g., in Switch Transformers [switch]. We further validate S-MBU accuracy on MoE models; detailed results are provided in Appendix A.4.
@@ 4.3 Benchmarking use cases of sparsity-aware performance metrics
Figure 6 illustrates the relationship between batch size and model sparsity, based on profiling using our custom benchmarking tool built on the latest vLLM release and evaluated on the MATH [MATH] dataset. In scenarios such as personal inference and post-training—characterized by small batch sizes (1–16)—all models exhibit high sparsity. At batch size 8, for instance, only 53.05%, 46.79%, and 18.44% of parameters are activated for DeepSeek-V2-Lite, Qwen1.5-MoE, and DeepSeek-R1, respectively. This significantly reduces bandwidth demands, making these models viable for deployment on cost-efficient hardware.
@@ 5 Benchmark Implementation
Dataset support. We evaluate all models on four representative benchmarks: MMLU, GSM8K, MATH, Arena-Hard, and LongBench. MMLU covers 57 diverse multiple-choice tasks to assess factuality and reasoning across domains. GSM8K and MATH focus on mathematical reasoning, with problems requiring multi-step solutions and short-form generation at varying difficulty levels. Arena-Hard evaluates long-form generation using 500 complex user queries from Chatbot Arena, judged by GPT-4-Turbo. It shows strong agreement with human preferences (89.1%) and better separability than other benchmarks. LongBench is a benchmark of long-context questions, making it well-suited for evaluating prefill-heavy workloads. Together, these benchmarks comprehensively test MoE LLMs across multiple-choice, short-form, and long-form generation tasks—covering knowledge, reasoning, and output quality—and are widely used in leading LLM leaderboards [open-llm-leaderboard, llm-perf-leaderboard, hallucinations-leaderboard].
@@ A.1 MoE-CAP Limitation and Future Work
Toward real-world deployment scenarios. We also aim to evaluate MoE systems in diverse cloud deployment settings, including serverless model endpoints [fu2024serverlessllm], elastic infrastructure [wagenlander2024tenplex, wang2023gemini, mai2020kungfu], and spot-instance pricing environments [mao2024skyserve, miao2024spotserve, wagenlander2020spotnik]. This will enable more realistic assessments of cost-performance trade-offs under variable resource conditions.
@@ A.7 Practical bandwidth and compute requirement calculation
We calculated theoretical performance through existing tools like LLM-Analysis [llm-analysis-chengli] and LLM-Viewer [yuan2024llm]. In practice, MoE systems often suffer inefficiencies due to redundant data transfers or calculations, resulting in performance losses. Consequently, the calculated theoretical bandwidth or OPS may not meet the required application performance (e.g., throughput). To determine the practical hardware requirements for a given set of inputs, it is essential to consider hardware utilization. Thus, the practical performance requirement is defined as: